A mixture of hydrogen and iodine gases are allowed to come to equilibrium at 445 degrees Celsius to form hydrogen iodide . Given the measured equilibrium concentrations of [H2]= 0.11 M, [I2] = 0.11 M and [HI] = 0.78 M, determine the value of the equilibrium constant .


1
Expert's answer
2021-02-09T04:42:43-0500

Equilibrium constant Kc = "\frac{[HI]^2}{[H_2][I_2]}"

Kc = "\frac{(0.78)^2}{(0.11)(0.11)}"


Kc = "\frac{0.6084}{0.0121}"


Kc = 50.28